OpenStudy (danielajohana):

simplify (1/x+1/y)/xy

OpenStudy (ipwnbunnies):

Use distributive property.

OpenStudy (danielajohana):

\[\frac{ \frac{ 1 }{ x }+\frac{ 1 }{ y } }{ xy }\]

OpenStudy (ipwnbunnies):

OHH, now I see

OpenStudy (ipwnbunnies):

Combine (Add) the numerator to get one fraction, and do a simple division

OpenStudy (ipwnbunnies):

You add the two fractions in the numerator by giving them the same denominator, which makes this problem simple and nice